In a copper wire of diameter 2mm, 3×1023 electrons flow in one second. Calculate the current density magnitude


1
Expert's answer
2022-03-01T17:33:03-0500

The current by definition is:


"I = \dfrac{ne}{t}"

where "n = 3\times 10^{23}" is the number of electrons, "e = 1.6\times 10^{-19}C" is the charge of the electron, and "t = 1s". Thus, obtain:


"I = 4.8\times 10^4A"

The charge density is:


"j = \dfrac{I}{\pi d^2/4}"

where "d = 2\times 10^{-3}m" is the diameter and "\pi d^2/4" is the cross-sectional area. Thus, obtain:


"j = \dfrac{4\cdot 4.8\times 10^4}{\pi \cdot (2\times 10^{-3})} \approx 153A/m^2"